# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ A.1 Proofs of the Auxiliary Lemmas and Resutls
$\|\bm{\Upsilon}_{1,j}-\bm{\Upsilon}_{2,j}\|^{2}_{F}$
$\quad\leq\|\bm{\Upsilon}_{1,j-1}-\bm{\Upsilon}_{2,j-1}\|^{2}_{F}+\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F},$ (A.7)
$\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}$
$\quad\leq\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{F}+\|\bm{U}_{1,j}\bm{V}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,j}\bm{V}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-1}^{-1}-\bm{C}_{2,j-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\|\bm{U}_{2,j}\|^{2}_{\mathrm{op}})$
$\quad\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}+\sum_{k=1}^{j}\big{[}\|\bm{U}_{1,k}\bm{V}_{1,k}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,k}\bm{V}_{2,k}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,k-1}^{-1}\|^{2}_{\mathrm{op}}$ (A.8)
$\quad+\|\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}\|\bm{V}_{1,k}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}\|\bm{U}_{2,k}\|^{2}_{\mathrm{op}})\big{]},$
$\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}\leq\bm{B}_{1}(1+\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}\max_{j}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\max_{j}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})^{j}+....$
$\sum_{j=0}^{p-1}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{F}$
$\quad\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}+\sum_{j=0}^{p-1}\sum_{k=1}^{j}\big{[}\|\bm{U}_{1,k}\bm{V}_{1,k}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,k}\bm{V}_{2,k}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,k-1}^{-1}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}\|\bm{V}_{1,k}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}\|\bm{U}_{2,k}\|^{2}_{\mathrm{op}})\big{]}$
$\quad\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}+\sum_{j=1}^{p-1}\sum_{k=1}^{j}\big{[}M_{1,U}M_{1,V}M_{2,U}M_{2,V}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}M_{1,V}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}M_{2,U})\big{]}$
$\quad\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{F}+p\sum_{k=1}^{p}\big{[}M_{1,U}M_{1,V}M_{2,U}M_{2,V}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{F}+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}$
$\quad+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}M_{1,V}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}M_{2,U})\big{]}.$
@@
$\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{F}$
$\quad\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{F}(1+\|\bm{C}_{1,j}^{-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})$
$\quad+2\|\bm{C}_{1,j}^{-1}\bm{U}_{1,j}\bm{V}_{1,j}-\bm{C}_{2,j}^{-1}\bm{U}_{2,j}\bm{V}_{2,j}\|^{2}_{F}$
$\quad\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{F}(1+\|\bm{C}_{1,j}^{-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})$
$\qquad+2\|\bm{C}^{-1}_{1,j}-\bm{C}^{-1}_{2,j}\|^{2}_{F}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}$
$\qquad+2\|\bm{C}_{2,j}^{-1}\|^{2}_{\mathrm{op}}\bigg{[}\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}\|\bm{V}_{2,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\bigg{]}$
$\quad\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{F}(1+\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{1,U}M_{1,V})$
$\qquad+2\|\bm{C}^{-1}_{1,j}-\bm{C}^{-1}_{2,j}\|^{2}_{F}M_{1,U}M_{1,V}$
$\qquad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\bigg{[}\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{F}M_{2,V}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{F}M_{1,U}\bigg{]},$
where $M_{\ell,U}=\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}},M_{\ell,V}=\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}$. Let $M_{\ell,C,U,V}=(1+\|\bm{C}_{\ell,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{\ell,U}M_{\ell,V})^{p}$, then
$\|\bm{\Upsilon}_{1,p-1}^{-1}\bm{\kappa}_{1,p}^{\mathrm{T}}-\bm{\Upsilon}_{2,p-1}^{-1}\bm{\kappa}_{2,p}^{\mathrm{T}}\|^{2}_{F}$
$\quad\leq 2M_{1,U}M_{1,V}M_{1,C,U,V}\sum_{k=0}^{p}\|\bm{C}^{-1}_{1,k}-\bm{C}^{-1}_{2,k}\|^{2}_{F}+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{F}$
$\qquad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{F}M_{2,U}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{F}.$
@@
We have $\bm{U}_{j}=\bm{C}_{j-1}\bm{L}_{j}(\bm{I}+\bm{L}^{\mathrm{T}}_{j}\bm{C}_{j-1}\bm{L}_{j})^{-1/2}$ and $\bm{V}_{j}=\bm{K}_{j}(\bm{K}_{j}^{\mathrm{T}}\bm{D}_{j-1}^{-1}\bm{K}_{j})^{-1/2}$.
$\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{j-1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j}\|^{2}_{\mathrm{op}}$ and $\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{D}_{j-1}^{-1}\|^{2}_{\mathrm{op}}$
@@
$\|\bm{V}_{j,1}-\bm{V}_{j,2}\|^{2}_{F}\leq r\|\bm{V}_{j,1}-\bm{V}_{j,2}\|^{2}_{\mathrm{op}}$
$\quad\leq\|\bm{K}_{1,1}\|^{2}_{\mathrm{op}}\|(\bm{K}_{j,1}^{\mathrm{T}}\bm{D}_{j-1,1}^{-1}\bm{K}_{1,1})^{-1}-(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{j-2,2}^{-1}\bm{K}_{1,2})^{-1}\|_{\mathrm{op}}$
$\qquad+r\|(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{0,2}^{-1}\bm{K}_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{K}_{j,1}-\bm{K}_{j,2}\|^{2}_{\mathrm{op}}$

### v2

@@ A.1 Proofs of the Auxiliary Lemmas
From the relation $\bm{D}_{j}=\bm{D}_{j-1}-\bm{W}_{j}^{\mathrm{T}}\bm{C}_{j-1}^{-1}\bm{W}_{j}$ with $\bm{W}_{j}=\bm{U}_{j}\bm{V}_{j}$.
$\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{\mathrm{F}}+\|\bm{W}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{W}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-1}^{-1}-\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{F}}$
$+(\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}})\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{\mathrm{F}}.$
We have the bounds
$\|\bm{W}_{1,j}-\bm{W}_{2,j}\|^{2}_{\mathrm{F}}\leq\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{U}_{2,j}\|^{2}_{\mathrm{op}},$
$\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}$, and $\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}$. Using the quadratic factorization $\bm{G}\bm{H}\bm{G}^{T}$ for $\bm{\Upsilon}^{-1}$ given in the proof of Lemma 2, we have the following,
$\|\bm{\Upsilon}^{-1}_{1,j}-\bm{\Upsilon}^{-1}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{G}_{1}-\bm{G}_{2}\|_{\mathrm{F}}+(\|\bm{G}_{1}\|^{2}_{\mathrm{op}}+\|\bm{G}_{2}\|^{2}_{\mathrm{op}})\|\bm{H}_{1}-\bm{H}_{2}\|^{2}_{\mathrm{F}}$
$\leq\|\bm{\Gamma}^{-1}_{1,j-1}-\bm{\Gamma}^{-1}_{2,j-1}\|^{2}_{\mathrm{F}}+\|\bm{D}^{-1}_{1,j}-\bm{D}^{-1}_{2,j}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}.$
Now,
$\|\bm{D}^{-1}_{1,j}-\bm{D}^{-1}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{D}^{-1}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{D}^{-1}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$
$\leq\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$
since $\|\bm{\Upsilon}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{\Upsilon}_{1,p}^{-1}\|^{2}_{\mathrm{op}},\|\bm{\Upsilon}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}$. Let
$\bm{P}={\begin{bmatrix}\bm{A}_{1}&\bm{A}_{2}\\ \bm{A}_{3}&\bm{A}_{4}\end{bmatrix}}^{-1}={\begin{bmatrix}\bm{B}_{1}&\bm{B}_{2}\\ \bm{B}_{3}&\bm{B}_{4},\end{bmatrix}}$
where
$\bm{B}_{1}$ $=(\bm{A}_{1}-\bm{A}_{2}\bm{A}_{4}^{-1}\bm{A}_{3})^{-1},$
$\bm{B}_{2}$ $=-(\bm{A}_{1}-\bm{A}_{2}\bm{A}_{4}^{-1}\bm{A}_{3})^{-1}\bm{A}_{2}\bm{A}_{4}^{-1},$
$\bm{B}_{3}$ $=-\bm{A}_{4}^{-1}\bm{A}_{3}(\bm{A}_{1}-\bm{A}_{2}\bm{A}_{4}^{-1}\bm{A}_{3})^{-1},$
$\bm{B}_{4}$ $=\bm{A}_{4}^{-1}+\bm{A}_{4}^{-1}\bm{A}_{3}(\bm{A}_{1}-\bm{A}_{2}\bm{A}_{4}^{-1}\bm{A}_{3})^{-1}\bm{A}_{2}\bm{A}_{4}^{-1}.$
Setting $\bm{P}=\bm{\Upsilon}_{j-1},\bm{A}_{1}=\bm{\Gamma}(0),\bm{A}_{2}=\bm{\xi}_{j-1}^{\mathrm{T}},\bm{A}_{3}=\bm{\xi}_{j-1},\bm{A}_{4}=\bm{\Upsilon}_{j-2}$, we have
$\bm{\Upsilon}_{j-1}^{-1}\bm{\kappa}_{j}^{\mathrm{T}}=\{\bm{C}_{j-1}^{-1}\bm{U}_{j}\bm{V}_{j},\bm{\Upsilon}_{j-2}^{-1}\bm{\kappa}_{j-1}^{\mathrm{T}}-\bm{\Upsilon}_{j-2}^{-1}\bm{\xi}_{j-1}^{\mathrm{T}}\bm{C}_{j-1}^{-1}\bm{U}_{j}\bm{V}_{j}\},$
since $\bm{\Gamma}(j)=\bm{U}_{j}\bm{V}_{j}^{\mathrm{T}}+\bm{\xi}_{j-1}\bm{\Gamma}^{-1}_{j-2}\bm{\kappa}_{j-1}$.
Now, $\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{op}},\|\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{op}}\leq 1$, and hence the relation
$\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\xi}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\xi}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}=\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$
holds. ∎
@@
We have the relations
$\|\bm{\Upsilon}_{1,j}-\bm{\Upsilon}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{\Upsilon}_{1,j-1}-\bm{\Upsilon}_{2,j-1}\|^{2}_{\mathrm{F}}+\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{\Upsilon}_{1}(0)\|^{2}_{\mathrm{op}}+\|\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{op}})\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}},$
$\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq\|\bm{D}_{1,j-1}-\bm{D}_{2,j-1}\|^{2}_{\mathrm{F}}$
$\quad+\|\bm{U}_{1,j}\bm{V}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,j}\bm{V}_{2,j}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,j-1}^{-1}-\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{C}_{1,j-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,j-1}^{-1}\|^{2}_{\mathrm{op}})$
×$\qquad\times(\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{U}_{2,j}\|^{2}_{\mathrm{op}})$
$\leq\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\quad+\sum_{k=1}^{j}\big{[}\|\bm{U}_{1,k}\bm{V}_{1,k}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,k}\bm{V}_{2,k}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{F}}$
$\quad+(\|\bm{C}_{1,k-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{op}})$
×$\qquad\times(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}\|\bm{V}_{1,k}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}\|\bm{U}_{2,k}\|^{2}_{\mathrm{op}})\big{]},$
$\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$ $\leq\bm{B}_{1}(1+\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}\max_{j}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\max_{j}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})^{j}+\cdots$
and hence
$\sum_{j=0}^{p-1}\|\bm{D}_{1,j}-\bm{D}_{2,j}\|^{2}_{\mathrm{F}}$ $\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\quad+\sum_{j=0}^{p-1}\sum_{k=1}^{j}\big{[}\|\bm{U}_{1,k}\bm{V}_{1,k}\|^{2}_{\mathrm{op}}\|\bm{U}_{2,k}\bm{V}_{2,k}\|^{2}_{\mathrm{op}}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{F}}$
$\qquad+(\|\bm{C}_{1,k-1}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{op}})(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}\|\bm{V}_{1,k}\|^{2}_{\mathrm{op}}$
$\qquad+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}\|\bm{U}_{2,k}\|^{2}_{\mathrm{op}})\big{]}$
$\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\quad+\sum_{j=1}^{p-1}\sum_{k=1}^{j}\big{[}M_{1,U}M_{1,V}M_{2,U}M_{2,V}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{F}}$
$\qquad+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})$
×$\qquad\qquad\times(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}M_{1,V}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}M_{2,U})\big{]}$
$\leq p\|\bm{\Upsilon}_{1}(0)-\bm{\Upsilon}_{2}(0)\|^{2}_{\mathrm{F}}$
$\quad+p\sum_{k=1}^{p}\big{[}M_{1,U}M_{1,V}M_{2,U}M_{2,V}\|\bm{C}_{1,k-1}^{-1}-\bm{C}_{2,k-1}^{-1}\|^{2}_{\mathrm{F}}$
$\qquad+(\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}+\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}})$
×$\qquad\qquad\times(\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}M_{1,V}+\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}M_{2,U})\big{]},$
completing the proof. ∎
Proof of Lemma 13.
We have
$\|\bm{\Upsilon}_{1,j-1}^{-1}\bm{\kappa}_{1,j}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-1}^{-1}\bm{\kappa}_{2,j}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$
$\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{F}}(1+\|\bm{C}_{1,j}^{-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})$
$\quad+2\|\bm{C}_{1,j}^{-1}\bm{U}_{1,j}\bm{V}_{1,j}-\bm{C}_{2,j}^{-1}\bm{U}_{2,j}\bm{V}_{2,j}\|^{2}_{\mathrm{F}}$
$\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{F}}(1+\|\bm{C}_{1,j}^{-1}\|^{2}_{\mathrm{op}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}})$
$\quad+2\|\bm{C}^{-1}_{1,j}-\bm{C}^{-1}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\|\bm{V}_{1,j}\|^{2}_{\mathrm{op}}$
$\quad+2\|\bm{C}_{2,j}^{-1}\|^{2}_{\mathrm{op}}\bigg{[}\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{V}_{2,j}\|^{2}_{\mathrm{op}}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}\|\bm{U}_{1,j}\|^{2}_{\mathrm{op}}\bigg{]}$
$\leq\|\bm{\Upsilon}_{1,j-2}^{-1}\bm{\kappa}_{1,j-1}^{\mathrm{T}}-\bm{\Upsilon}_{2,j-2}^{-1}\bm{\kappa}_{2,j-1}^{\mathrm{T}}\|^{2}_{\mathrm{F}}(1+\|\bm{C}_{1,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{1,U}M_{1,V})$
$\quad+2\|\bm{C}^{-1}_{1,j}-\bm{C}^{-1}_{2,j}\|^{2}_{\mathrm{F}}M_{1,U}M_{1,V}$
$\qquad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\big{[}\|\bm{U}_{1,j}-\bm{U}_{2,j}\|^{2}_{\mathrm{F}}M_{2,V}+\|\bm{V}_{1,j}-\bm{V}_{2,j}\|^{2}_{\mathrm{F}}M_{1,U}\big{]},$
where $M_{\ell,U}=\max_{j}\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}$, and $M_{\ell,V}=\max_{j}\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}$.
Let $M_{\ell,C,U,V}=(1+\|\bm{C}_{\ell,p}^{-1}\|^{2}_{\mathrm{op}}\|M_{\ell,U}M_{\ell,V})^{p}$, then
$\|\bm{\Upsilon}_{1,p-1}^{-1}\bm{\kappa}_{1,p}^{\mathrm{T}}-\bm{\Upsilon}_{2,p-1}^{-1}\bm{\kappa}_{2,p}^{\mathrm{T}}\|^{2}_{\mathrm{F}}$ $\leq 2M_{1,U}M_{1,V}M_{1,C,U,V}\sum_{k=0}^{p}\|\bm{C}^{-1}_{1,k}-\bm{C}^{-1}_{2,k}\|^{2}_{\mathrm{F}}$
$\quad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}M_{2,V}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{U}_{1,k}-\bm{U}_{2,k}\|^{2}_{\mathrm{F}}$
$\quad+2\|\bm{C}_{2,p}^{-1}\|^{2}_{\mathrm{op}}\|^{2}_{\mathrm{F}}M_{2,U}M_{1,V}M_{1,C,U,V}\sum_{k=1}^{p}\|\bm{V}_{1,k}-\bm{V}_{2,k}\|^{2}_{\mathrm{F}}.$
@@
We have
$\bm{U}_{j}=\bm{C}_{j-1}\bm{L}_{j}(\bm{I}+\bm{L}^{\mathrm{T}}_{j}\bm{C}_{j-1}\bm{L}_{j})^{-1/2},\quad\bm{V}_{j}=\bm{K}_{j}(\bm{K}_{j}^{\mathrm{T}}\bm{D}_{j-1}^{-1}\bm{K}_{j})^{-1/2},$
and the bounds $\|\bm{U}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{C}_{j-1}\|^{2}_{\mathrm{op}}\|\bm{L}_{j}\|^{2}_{\mathrm{op}}$ and $\|\bm{V}_{\ell,j}\|^{2}_{\mathrm{op}}\leq\|\bm{D}_{j-1}^{-1}\|^{2}_{\mathrm{op}}$.
@@
$\|\bm{V}_{j,1}-\bm{V}_{j,2}\|^{2}_{\mathrm{F}}$ $\leq r\|\bm{V}_{j,1}-\bm{V}_{j,2}\|^{2}_{\mathrm{op}}$
$\leq\|\bm{K}_{1,1}\|^{2}_{\mathrm{op}}\|(\bm{K}_{j,1}^{\mathrm{T}}\bm{D}_{j-1,1}^{-1}\bm{K}_{1,1})^{-1}-(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{j-2,2}^{-1}\bm{K}_{1,2})^{-1}\|_{\mathrm{op}}$
$\quad+r\|(\bm{K}_{j,2}^{\mathrm{T}}\bm{D}_{0,2}^{-1}\bm{K}_{j,2})^{-1/2}\|^{2}_{\mathrm{op}}\|\bm{K}_{j,1}-\bm{K}_{j,2}\|^{2}_{\mathrm{op}}.$